OpenStudy (anonymous):

For each ordered pair, determine whether it is a solution to the system of equations. 9x-2y=8 -7x+3y=1 (2,5),(-8,1),(-4,-9),(0,-4)

jhonyy9 (jhonyy9):

so just substitute these numbers from parentheses on place of x and y and check it than are or not solution of the system ok ?

jhonyy9 (jhonyy9):

for example the first parantheses (2,5) x=2 y=5 9x-2y=8 -7x+3y=1 9*2 -2*5=8 -7*2+3*5=1 18-10=8 yes is true -14+15=1 yes is true so than this is right continue and try checking the second pair and ... so